$177,000 $238,000 $311,000 Initial investment (CF) Year (t) $82,000 68,000 55,000 58,000 60,000 Cash inflows (CF) $53,000 70,Risk classes and RADR Moses Manufacturing is attempting to select the best of three mutually exclusive projects, X, Y, and Z.
Answer #1

NPV is calculated by summing up all the cashflow from the project by discounting each cashflow by appropriate discount rate. in our case the discount rate is RADR.

NPV = - initial investment + CF1/(1+r) +CF2/(1+r)2 + CF3/(1+r)3 +CF4/(1+r)4 +CF5/(1+r)5

RADR for project X = 22.3%

RADR for project Y = 13.5%

RADR for project Z = 15.4%

Risk Adjusted NPV for project X = - 177000 + 82000/(1+22.3%) +68000/(1+22.3%)2 + 55000/(1+22.3%)3 +58000/(1+22.3%)4 + 60000/(1+22.3%)5 = $13,431.76

Risk Adjusted NPV for project Y = - 238000+ 53000/(1+13.5%) +70000/(1+13.5%)2 + 72000/(1+13.5%)3 +88000/(1+13.5%)4 + 94000/(1+13.5%)5 = $15,210.20

Risk Adjusted NPV for project Z = - 177000 + 94000/(1+15.4%) +94000/(1+15.4%)2 + 94000/(1+15.4%)3 +94000/(1+15.4%)4 + 94000/(1+15.4%)5 = $1,141.23

Since the risk adjusted NPV for Y is the highest, firm must take project Y.